Comsider the example of the number of languages spoken by Filipino school children. Define x to be the number of languages in which a randomly chosen Filipino child attending school can hold an everyday conversation. Assume the probability function of X, P(x), is as shown in the following table.


X - 1 2 3 4 5 6

P(x) - 0.663, 0.226, 0.066, 0.019, 0.004


a. What is the mean of X?

b. Find the variance of X.

c.Find the standard deviation of X.


1
Expert's answer
2022-02-01T12:53:20-0500

Solution:

X - 1 2 3 4 5 6

P(x) - 0.663, 0.226, 0.066, 0.019, 0.004, 0.022

(a)

"Mean=E(X)=\Sigma x.P(x)
\\=1\times 0.663+ 2\times0.226+ 3\times0.066+4\times 0.019+ 5\times0.004+6\times0.022
\\=1.541"

(b)

"E(X^2)=x^2.P(x)
\\=1^2\times 0.663+ 2^2\times0.226+ 3^2\times0.066+4^2\times 0.019+ 5^2\times0.004+6^2\times0.022
\\=3.357
\\Variance=\sigma^2=E(X^2)-[E(X)]^2
\\=3.357-1.541^2
\\=0.982319"

(c)

Standard deviation"=\sigma=\sqrt{0.982319}=0.99112007345"